# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

### v1

@@ IV Hierarchical Prompt Tuning (HiPTune)
Figure 3: (a) The VP-Tree establishes a hierarchical prompt tree for all attack types, organized into three levels of grouping categories. It assigns a learnable prompt for each attack type and integrates the hierarchical prompts into the final output. (b) The APP employs a hierarchical prompt pruning mechanism to select the most promising prompt for each sample. (c) The DPI aims to aggregate all candidate prompts from different levels of VP-Tree and projects them into the text encoder for subsequent similarity computation.
IV-A Preliminaries
@@
Vision-Language Models (VLM) like CLIP [54] utilize both text encoder $\mathcal{T}(\cdot)$ (like Transformer [55]) and image encoder $\mathcal{I}(\cdot)$ (like ViT [56]), where the input is the (text, image) pair, to leverages broader supervision. The outputs of both encoders are then dot-multiplied to obtain the probability score $p$:
$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\bm{w}_{\bm{i}},\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\bm{w}_{\bm{j}},\bm{f}\right)/\tau\right)},$ (1)
@@
where $\left\{\bm{w}_{i}\right\}_{i=1}^{K}$ denotes the set of weight vectors generated by the text encoder, and $\bm{f}$ represents the features extracted by the image encoder. $K$ is the number of classes, and $\tau$ refers to the temperature parameter learned by CLIP, $\cos(\cdot,\cdot)$ means the cosine similarity. Moreover, CoOp [57] explores prompt engineering in the VLM, constructing learnable prompts $\bm{t}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{M}[\mathrm{CLASS}]$ and feeding to the fixed Text Encoder. The $[\mathrm{V}]_{m}(m\in\{1,\ldots,M\})$ is the vectors with the same dimension as word embeddings, and $M$ is a hyperparameter specifying the number of context tokens. The prediction probability is further computed as follows:
$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{i}\right),\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{j}\right),\bm{f}\right)/\tau\right)}.$ (2)
@@
Visual Prompt Tuning (VPT) aims to fine-tune the pre-trained models with only a few learnable parameters. Given the input image tokens $\bm{x}_{i},i\in{1,...,l}$ with length $l$, the trainable vectors $\bm{P}=\left\{\bm{p}^{k}\in\mathbb{R}^{d}\mid k\in\mathbb{N},1\leq k\leq p\right\}$, where d is the dimension of the patch embeddings and $p$ is the number of prompts, are cat with the original tokens and fed into the fixed encoders $Input=cat[\bm{x}\cdot\bm{P}]$.
IV-B Overview of HiPTune
@@
We proposed a HiPTune that adopts hierarchical prompts to find multiple classification criteria from different semantic spaces and further integrates them, aiming to form more robust classification criteria for each input sample. Shown in Fig. 3, first, we construct the Visual Prompt Tree to gain the various semantics spaces tailored for each attack detection task. Then, given the input image $\bm{x}$, we adaptively select the most suitable prompt from each level of the VP-Tree by our Adaptive Prompt Pruning module. In addition, we integrate the prompt candidates from each selected level of the APP to describe the complete features of the input sample and feed it to the text encoder for better alignment.
IV-C Visual Prompt Tree (VP-Tree)
@@
Define the levels. We can obtain a hierarchical relationship between each attack type based on the spoof and forgery types. Apart from the subjective categorization from the dataset, the features of different attacks demonstrate a more compact grouping relative to other categories while maintaining distinct independence within each category. For a VP-Tree with $N$ degrees, the level ${Tree}^{i}(i\in{1,2,...,N})$ has $n$ prompts, and $n$ may vary according to the set-ups. In our case, we define $3$-levels for the VP-Tree shown as Fig. 3 (a). The first level distributes the most obvious data as live, physical, and digital attack subtrees. The physical subtree is first separated into 2D and 3D attacks, which represents the second level of the VP-Tree. Print, replay, and cutouts are three key types for the level 3 prompts of a 2D attack. As for the 3D attacks, the masks’ material is always considered to be the spoof clues, so we define transparent, plaster, and resin prompts within the 3D attack prompt. Concerning the digital subtree, we categorize the digital attacks into $3$ kinds according to their forgery techniques. The digital manipulation method includes attribute-editing, face-swap, and video-driven faces. Adversarial attacks are divided into pixel-level and semantic-level attacks for forgery areas. We also define $3$ sub-prompts for the digital generation attacks: ID-consistent generation, style transfer, and prompt-based editing forgeries.
@@
$\bm{p}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{l},$ (3)
@@
Train and integrate the prompts. In order to force each prompt to learn the classification criteria specific to each attack type, we train the prompts in a supervised manner during the first stage. Given the input $\bm{x}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel, height, and width of the input images, we manually record the attack types and select the corresponding prompts at each level accordingly. For the live face images, since we train our model with (live, fake) pairs, we choose the same prompt as the partner. Then, we average the selected prompts to form the final prompt and link with the image tokens. The final input of the image encoder is formulated as Eq. 4:
${\mathcal{I}}_{input}=cat[(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\cdot\bm{p}],$ (4)
@@
where $(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\in\mathbb{R}^{l+1,dim}$ is the image tokens after the patch embeddings. $l$ represents the length of image tokens, and $dim=768$ is the embedding dimension. The $\bm{p}$ is the final output of the integrated prompts as Eq. 5
$\bm{p}=mean(\bm{p}_{1}+\bm{p}_{2}+\bm{p}_{3}),$ (5)
where $\bm{p}_{1},\bm{p}_{2},\bm{p}_{3}$ are the selected prompts at level $1,2,3$ and $mean(\cdot)$ represents the average of the inputs.
@@
After fully training the VP-Tree to represent the classification criteria for multiple semantic spaces, the Adaptive Prompt Pruning (APP) module will learn to select the most suitable prompts for each sample.
IV-D Adaptive Prompt Pruning (APP)
@@
As shown Fig. 3 (b), aiming to explore classification criteria from different semantic spaces, the APP module adopts a hierarchical pruning mechanism to select the most suitable prompt for each sample. Given the input image $\bm{x}\in\mathbb{R}^{C\times H\times W}$, the patch embedding first projects the image into the visual embedding dimension and reshapes $\bm{x}$ into a series of image token with length $L$: $\bm{x}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{L\times D}$. In our case, $C=3$ represents the channels for the image, and $H,W$ are both $244$ for the input image, the token’s length $L=196$, and the dimension of visual embedding is $D=768$. After adding the positional embeddings, a class token is linked with the image token, and then the image token $\bm{x}\in\mathbb{R}^{197\times 768}$ is fed to the APP.
Level-1 prompt pruning. Given the significant distribution variance between physical and digital attacks, it is simple to determine whether physical or digital criteria should be used to classify this sample. Thus, we utilize a simple linear layer to select the level 1 prompt:
$G_{1}(\bm{x})=\mathbf{W}_{1}\cdot\bm{x}+bias_{1},$ (6)
where $G_{1}$ is the gate function of the level-1 prompt pruning and $\mathbf{W}_{1},bias_{1}$ are learnable weights and bias. The level-1 router first projects each patch’s features into three alternatives using $G_{1}$ and then calculates the mean value over all tokens. After applying Softmax between three classes, we choose the $TOP_{K=1}$ candidate prompt. The whole procedure can be described as:
$i^{p}_{1}=TOP_{k=1}(softmax(mean(G_{1}(\bm{x}),197))),$ (7)
where $i^{p}_{1}$ denotes the index of the selected prompt, $mean(\cdot,\cdot)$ represents the mean calculus over specific dimension. Furthermore, the selected prompt is linked with the original image tokens and forwarded to the next level.
Level-2 prompt pruning. Regarding the level-2 prompt pruning, it is challenging to separate samples from attack types within each category if we only rely on fixed visual embeddings. In light of this, we apply Central Difference Convolution (CDC) [58] and the Frequency-Adaptive Dilated Convolution (FADC) [59] for level-2 prompt pruning in the physical and digital branches, respectively.
After the CDC or the FADC, the output feature map is sent to a linear layer, and then the tokens’ scores are averaged to select the candidate prompt further. The level-2 gate functions can be expressed as:
$G_{2}^{0}$ $=\mathbf{W}_{2}^{0}\cdot CDC_{2}^{0}(\bm{x})+bias_{2}^{0}$ (8)
$G_{2}^{1}$ $=\mathbf{W}_{2}^{1}\cdot FADC_{2}^{1}(\bm{x})+bias_{2}^{1},$
and the selected prompt can be denoted as:
$i^{p}_{2}=TOP_{k=1}(softmax(mean(G_{2}(\bm{x}),197))).$ (9)
The level-1 and level-2 prompts are combined averagely and concat with the image tokens and fed to the next router.
Level-3 prompt pruning. The level-3 prompt pruning procedure is the same as the level-2 pruning process. The gate functions can be described as follows:
$G_{3}^{i}=\mathbf{W}_{3}^{i}\cdot Conv_{3}^{i}(\bm{x})+bias_{3}^{i},$ (10)
where $i\in{1,2,3}$ for physical 2D/3D, digital manipulation, generation, and $i\in{1,2}$ for digital adversarial attacks.
Finally, the selected prompts from all three levels are averaged together and concated with the original image tokens. For the live images, the live prompt from the first level will be directly concated with the tokens and fed to the image encoder.
IV-E Dynamic Prompt Interaction (DPI)
@@
Specifically, first, we directly pass the live prompts from the first level as the live class text prompts. As for the fake prompts, we apply the probabilities that the app has predicted to the candidate prompts and combine them weightedly. For the first level, we apply the softmax function only to physical and digital prompt scores to obtain the binary probabilities of two categories. Then, we weighted sum the prompts to get a weighted level-1 fake prompt. In the second level, due to the router from the first level choosing a subtree for the sample and needing a possible fake prompt specific to the input sample, we repeat the weighted sum of candidate prompts until the router has not predicted the further probabilities. Lastly, we form prompts for both live and fake faces and send them to the text encoder for further similarity computation.
IV-F Objectives and Inference
The training objective for the first training stage is to help the prompts to learn the semantic information and classification criteria specific to the attack type. The second stage aims to train the APP to select the most suitable prompts tailored for each sample successfully. The loss for both stages is:
$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{triple},$ (11)
where $\mathcal{L}_{CE}$ represents the cross-entropy loss and $\mathcal{L}_{triple}$ is the Asymmetric Triplet Loss [60]. The reason for applying Asymmetric Triplet Loss is to separate the different prompts to preserve their unique classification criteria while combat the relatively similar prompts.
V Experiments
V-A Experiment Settings
Datasets, Protocols, and Evaluation Metrics. We evaluated the effectiveness of the proposed UniAttackData+ dataset and the robustness of the HiPTune algorithm using three existing UAD datasets. For JFSFDB, we conducted both internal and external testing within its original FAS and FFD datasets, as well as within the joint training dataset. For UniAttackData, we tested using its original p1, p2.1, and p2.2 subsets. For our proposed UniAttackData+, we performed training and testing based on the four protocols we introduced. Additionally, we established a UAD-Benchmark to assess the model’s performance in cross-dataset testing. We evaluate model performance with four metrics: (1) ACER (average of FRR and FAR), (2) AUC (theoretical model performance), (3) EER (optimal threshold), and (4) ACC (overall accuracy).

### v3

@@ 4 Hierarchical Prompt Tuning (HiPTune)
Figure 4: (a) The VP-Tree establishes a hierarchical prompt tree for all attack types, organized into three levels of grouping categories. It assigns a learnable prompt for each attack type and integrates the hierarchical prompts into different levels of the encoder. (b) The DPI aims to aggregate all candidate prompts from different levels of the VP-Tree and projects them into the text encoder for subsequent similarity computation. (c) The APP employs an adaptive prompt pruning mechanism to select the most promising prompt for each sample.
4.1 Preliminaries
@@
Vision-Language Models (VLMs) like CLIP [83] utilize both text encoder $\mathcal{T}(\cdot)$ (like Transformer [84]) and image encoder $\mathcal{I}(\cdot)$ (like ViT [85]), where the input is the image-text pair, providing a wide range of semantic information as supervision. The outputs of both encoders are then dot-multiplied to obtain the probability score $p$:
$p(y=i\mid\boldsymbol{x})=\frac{\exp\left(\cos\left(\boldsymbol{w}_{\boldsymbol{i}},\boldsymbol{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\boldsymbol{w}_{\boldsymbol{j}},\boldsymbol{f}\right)/\tau\right)},$ (1)
@@
where $\left\{\boldsymbol{w}_{i}\right\}_{i=1}^{K}$ denotes the set of weight vectors generated by the text encoder, and $\boldsymbol{f}$ represents the features extracted by the image encoder. $K$ is the number of classes, and $\tau$ refers to the temperature parameter learned by CLIP, $\cos(\cdot,\cdot)$ means the cosine similarity. Moreover, CoOp [86] explores prompt engineering in the CLIP, constructing learnable prompts $\boldsymbol{t}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{M}[\mathrm{CLASS}]$ and feeding to the frozen Text Encoder. The $[\mathrm{V}]_{m}(m\in\{1,\ldots,M\})$ is the vectors with the same dimension as word embeddings, and $M$ is a hyperparameter specifying the number of context tokens. The prediction probability is further computed as follows:
$p(y=i\mid\boldsymbol{x})=\frac{\exp\left(\cos\left(\mathcal{T}\left(\boldsymbol{t}_{i}\right),\boldsymbol{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\mathcal{T}\left(\boldsymbol{t}_{j}\right),\boldsymbol{f}\right)/\tau\right)}.$ (2)
@@
Visual Prompt Tuning (VPT) aims to fine-tune the pre-trained models with only a few learnable parameters. Given the input image tokens $\boldsymbol{x}_{i},i\in{1,...,l}$ with length $l$, the trainable vectors $\boldsymbol{P}=\left\{\boldsymbol{p}^{k}\in\mathbb{R}^{d}\mid k\in\mathbb{N},1\leq k\leq p\right\}$, where $d$ is the dimension of the patch embeddings and $p$ is the number of prompts, are cat with the original tokens and fed into the frozen encoders $Input=cat[\boldsymbol{x}\cdot\boldsymbol{P}]$ in certain orders.
4.2 Overview of HiPTune
@@
We proposed a HiPTune that adopts hierarchical prompts to find multiple classification criteria from different semantic spaces and further integrates them, aiming to form more robust classification criteria for each input sample. Shown in Fig. 4, first, we construct the Visual Prompt Tree (VP-Tree) to gain the various semantics spaces tailored for each attack detection task. Then, given the input image $\boldsymbol{x}$, we adaptively select the most suitable prompt from each level of the VP-Tree by our Adaptive Prompt Pruning (APP) module. In addition, we integrate the prompt candidates from each selected level of the APP to describe the complete features of the input sample and feed it to the text encoder for better alignment through our Dynamic Prompt Integration (DPI) mechanism.
4.3 Visual Prompt Tree (VP-Tree)
Inspired by the work of [87, 88], learnable prompt parameters can be used to adapt a pretrained encoder, steering its output features to better align with downstream tasks. In the context of the UAD task, the diversity and high realism of forgery techniques pose significant challenges. To address this, we propose a hierarchical prompt tree designed to deliver coarse-to-fine guidance throughout the layered processing of a Transformer. At each block, the prompts encourage the encoder to refine its feature representations progressively, improving discrimination for classification. Notably, our prompts are class-agnostic in the sense that they target the distinction between authentic and forged samples in general, applying a specific prompt to each “real–fake” sample pair to capture their distinguishing characteristics.
@@ 4.3 Visual Prompt Tree (VP-Tree)
Define the levels. We can obtain a hierarchical relationship between each attack type based on the spoof and forgery types from Fig. 4. Apart from the subjective categorization from the dataset, the features of different attacks demonstrate a more compact grouping relative to other categories while maintaining distinct independence within each category. For a VP-Tree with $N$ degrees, the level ${Tree}^{i}(i\in{1,2,...,N})$ has $n$ prompts, and $n$ may vary according to the set-ups. In our case, we define $3$-levels for the VP-Tree shown as Fig. 4 (a). The first level distributes the most obvious data as physical and digital attack subtrees. The physical subtree is first separated into 2D and 3D attacks, which represent the second level of the VP-Tree. Print, replay, and cutouts are three key types for the level 3 prompts of a 2D attack. As for the 3D attacks, the masks’ material is always considered to be the spoof clues, so we define transparent, plaster, and resin prompts within the 3D attack prompt. Concerning the digital subtree, we categorize the digital attacks into $3$ kinds according to their forgery techniques. The digital manipulation method includes attribute editing, face swapping, and video-driven faces. Adversarial attacks are divided into pixel-level and semantic-level attacks for forgery areas. We also define $3$ sub-prompts for the digital generation attacks: ID-consistent generation, style transfer, and prompt-based editing forgeries.
It is worth noting that to ensure continuity in the properties and semantics at each level, each subsequent prompt partially inherits semantic information from its preceding prompt. Specifically, given a prompt defined in $\boldsymbol{p}^{0}\in\mathbb{R}^{l\times D}$, we apply an inheritance coefficient $\alpha$ to generate the inherited prompt $\boldsymbol{p}^{1}=cat[\boldsymbol{p}^{1}\in\mathbb{R}^{(l-\alpha l)\times D},\boldsymbol{p}^{1}\in\mathbb{R}^{\alpha l\times D}]$. This inherited component is then proportionally compressed, every $\frac{\alpha}{1}\cdot l$ tokens have been calculated the mean value and concatenated to the next-level prompt. This design ensures that the new prompt retains essential characteristics from its parent while introducing its own fine-grained, task-specific distinctions.
@@
$\boldsymbol{p}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{l},$ (3)
@@
Train and integrate the prompts. In order to force each prompt to learn the classification criteria specific to each attack type, we train the prompts in a supervised manner during the first stage. Given the input $\boldsymbol{x}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel, height, and width of the input images, we manually record the attack types and select the corresponding prompts at each level accordingly. Following the approach of XX, we insert the prompt tokens between the original image tokens and the image class token. This design allows the prompts to effectively modulate the flow of information from the input representations to the classification token, encouraging the encoder to integrate the prompt’s guidance during feature aggregation and thereby refining task-relevant representations. The final input of each Transformer block is formulated as Eq. 4:
${\mathcal{I}}_{input}=cat[\boldsymbol{x}_{cls},\boldsymbol{p},(\boldsymbol{x}_{1},...,\boldsymbol{x}_{l})]$ (4)
@@
where $(\boldsymbol{x}_{cls},\boldsymbol{x}_{1},...,\boldsymbol{x}_{l})\in\mathbb{R}^{l+1,dim}$ is the image tokens after the patch embeddings. $l$ represents the length of image tokens, and $dim=768$ is the embedding dimension. The $\boldsymbol{p}$ is the selected prompt for each level.
As illustrated in the figure, we integrate different levels of prompts into specific layers of the Transformer blocks. As the Transformer progressively deepens, the semantic representation of tokens becomes increasingly precise. To leverage this property, we insert the first-level prompt after $N_{0}$ blocks—once the features have developed sufficient basic representational capacity—focusing on coarse discrimination between physical and digital attacks. Subsequently, after an additional $N_{1}$ and $N_{2}$ blocks, we introduce the second and third-level prompts, which provide increasingly fine-grained semantic guidance. Finally, the features undergo further processing through $N_{3}$ additional blocks to yield the final representation for classification.
@@
After fully training, the VP-Tree to represent the classification criteria for multiple semantic spaces, the Adaptive Prompt Pruning (APP) module will learn to select the most suitable prompts for each sample.
TABLE IV: Results on UniAttackData+. CLIP-V [83] indicates removing the text encoder from CLIP. CoOp [86] (unified/specific) represents the CoOp without/with the class-specific context.
Model P1: ID P2: cross-attack-methods P3.: cross-attacks-types
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V [83] 9.36 92.46 88.05 9.36 28.77 71.99 78.49 28.77 49.87 43.12 58.08 49.87
CLIP [83] 9.93 95.22 83.75 9.93 22.89 72.79 76.15 22.89 34.37 58.09 59.90 34.37
CoOp [86] 14.76 90.89 83.92 14.76 43.14 59.64 47.83 43.14 72.00 20.89 27.18 72.00
CDCN++ [89] 12.77 94.53 90.89 11.72 15.23 91.78 84.72 15.22 37.55 67.90 58.67 38.46
CFPL-FAS [90] 13.93 93.77 96.24 11.65 16.10 86.34 83.09 16.10 29.60 73.50 71.46 29.60
MoE-FFD [91] 17.30 91.90 96.50 15.69 30.53 70.09 92.82 31.39 46.01 51.27 89.14 47.97
LLA-Net [92] 10.57 95.64 95.97 7.61 18.34 82.97 89.86 21.31 36.58 57.02 68.25 42.20
FA3-CLIP [59] 8.26 93.96 94.53 8.45 35.81 63.89 88.51 47.37 44.44 48.34 56.95 41.05
HiPTune 7.31 97.87 90.02 7.61 14.26 92.50 90.26 14.71 25.68 77.62 69.34 26.02
4.4 Adaptive Prompt Pruning (APP)
@@
As shown in Fig. 4 (a) and (c), aiming to explore classification criteria from different semantic spaces, the APP module adopts a hierarchical pruning mechanism to select the most suitable prompt for each sample. Given the input image $\boldsymbol{x}\in\mathbb{R}^{C\times H\times W}$, the patch embedding first projects the image into the visual embedding dimension and reshapes $\boldsymbol{x}$ into a series of image tokens with length $L$: $\boldsymbol{x}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{L\times D}$. In our case, $C=3$ represents the channels for the image, and $H,W$ are both $224$ for the input image, the token’s length $L=196$, and the dimension of visual embedding is $D=768$. After adding the positional embeddings, a class token is linked with the image token, and then the image token $\boldsymbol{x}\in\mathbb{R}^{197\times 768}$ is fed to the Transformer blocks. Through $N_{0}$ blocks, the image tokens and the level-1 prompts are fed into the APP.
To select the most appropriate prompt, we first designate the front-most element (the first token after concatenation, with length 1) as the representative of each prompt. For the input image tokens, we use their $[CLS]$ token as the representation of the entire image. We then compute the attention scores between all prompt representatives and the image $[CLS]$ token. After applying the softmax function to normalize these scores, we select the single most relevant prompt using a $TOP_{k=1}$ strategy. This selected prompt is then inserted back into the sequence of image tokens to guide further encoding. The overall procedure is formally described as follows:
$index=TOP_{k=1}[Attention(\textbf{p}[0,:]\cdot x[0,:,:])],$ (5)
where $Attention(\cdot)$ represents the attention calculation, and the $index$ is the selection index of the most suitable prompt.
It is important to note that to ensure each prompt accurately captures its intended semantic information during selection, we supervise the training of prompts using ground-truth labels at each level for different samples in the training set. This enables each prompt to specialize in discriminating its designated category or property. Once trained, the resulting VP-Tree is integrated into the model as a fixed module, supporting the APP (Adaptive Prompt Pooling) selection mechanism during downstream training.
4.5 Dynamic Prompt Interaction (DPI)
@@
Specifically, shown as Fig. 4 (b), after the APP module selects the appropriate prompts at each hierarchical level, we stack these prompts in order to form a combined sequence. This stacked sequence is then processed using self-attention, allowing information exchange across levels. Following the softmax operation, we compute a weighted sum of the three hierarchical prompts to produce the final unified prompt. Finally, this prompt is projected from the original image feature dimension to the target text feature dimension through a dedicated mapping layer.
5 Experiments
5.1 Experiment Settings
Datasets, Protocols, and Evaluation Metrics. We evaluate the effectiveness of the proposed UniAttackData+ and HiPTune using three existing UAD datasets. For the UniAttackData+ dataset, we conduct experiments on the three proposed protocols. For JFSFDB [38] dataset, we only conduct cross-domain experiments on both separate training on each single task (i.e., PAD or FFD) and joint training within both two tasks (i.e., UAD). Specifically, separate training is conducted by training two distinct models on datasets SiW, 3DMAD, HKBU for PAD, and FF++ for FFD, respectively. The trained models are then evaluated on cross-domain datasets MSU, 3DMask, ROSE for PAD and DFDC, CelebDFv2 for FFD. Instead of training separate models, joint training involves learning a unified model for both tasks by leveraging all data from SiW, 3DMAD, HKBU, FF++. The unified model is then evaluated on cross-domain datasets MSU, 3DMask, ROSE, DFDC, CelebDFv2 for UAD. For UniAttackData [39] dataset: (1) Protocol 1 evaluates performance under a unified attack detection task with all attack types included in training and testing; (2) Protocol 2 assesses generalization to unseen attack types using a leave-one-type-out strategy, with Protocols 2.1 and 2.2 focusing on unseen physical and digital attacks, respectively. Following the setup in FA3-CLIP [59], we also conduct experiments under Protocol 1.1, 1.2, and 1.3. Protocols 1.1 and 1.2 independently evaluate the digital category by excluding deepfake and adversarial attacks, respectively, from training and validation, testing them only on separate identity sets. Protocol 1.3 includes all digital subtypes with standard distribution. These protocols assess the model’s robustness in unified face attack detection across both physical and digital threats. In addition, we construct a domain generalization-based UAD benchmark using the JFSFDB, UniAttackData, and UniAttackData+ datasets, where any two are combined for training and the remaining one is used for testing.
We evaluate performance using four metrics: (1) ACER, the average of the false rejection rate (FRR) and the false acceptance rate (FAR), (2) AUC, reflecting the model’s theoretical performance, (3) EER, representing the error rate at the optimal threshold, and (4) ACC, indicating the overall classification accuracy.